A 500 kg cannon rests on wheels that are free to roll. The cannon fires a 2 kg cannon ball horizontally with an average force of 400 N. What is the acceleration of the cannon ball? What is the recoil acceleration of the cannon
1
Expert's answer
2012-11-14T07:11:52-0500
acceleration is a=f/m
f=400N m=2 kg
a=200 m/s^2
recoil acceleration due to II Newtons law is
m1a1=F
a1=F/m1
a1=400/200= 2 m/s^2